The length of a rectangle is 8 inches more than its width. The perimeter of the rectangle is 24 inches. What are the width and length of the rectangle?
Answer by checkley77(12844)

L=2W+8
2L+2W=24
2(2W+8)+2W=24
4W+16+2W=24
6W=24-16
6W=8
W=8/6
W=4/3 ANS.
L=2*4/3+8
L=8/3+8
L=8/3+24/3
L=32/3 ANS.
PROOF:
2*4/3+2*32/3=24
8/3+64/3=24
72/3=24
24=24
2+20